# On the hierarchical Bayesian modelling of frequency response functions

### v1

@@ 3 Novelty and Contribution
Unlike previous efforts, the current work is concerned with developing a probabilistic FRF model, using an FRF equation based on modal parameters (natural frequency, mode shape, and damping) as the likelihood function. As with L.A. Bull et al. (2022, 2023a, 2023b), a hierarchical Bayesian (or MTL) framework was used. However, the focus is data sharing in the presence of low-resolution frequency information (i.e., given sparse data). The inability to localise both time and frequency to a fine resolution, per the uncertainty principle, means that missing time-domain data (e.g., acceleration, velocity) measured from SHM sensors will result in fewer spectral lines in the frequency domain. This decreased frequency resolution could impair proper characterisation/identification of modal peaks, which may be features of interest in a damage-detection application. By using a combined-inference approach, structures whose FRFs have many spectral lines in a band of interest can lend statistical support to those whose FRFs are limited by missing data. In addition, the current work incorporated functional relationships to describe changes in environmental conditions. This inclusion of functional relationships allows for extrapolation to temperature states not used in model training, which increases the amount of normal-condition information available, to address data-scarcity challenges. In contrast to L.A. Bull et al. (2023b), parameters were learnt over the experimental campaign, rather than hyperparameters, giving greater physical interpretability of the results.
@@ 4 Background theory
In this work, combined probabilistic FRF models were developed using a hierarchical Bayesian modelling approach, to support PBSHM research. This section provides an overview of the methodologies used, including hierarchical multilevel modelling from a Bayesian perspective, linear modal analysis, FRF estimation, and model evaluation.
@@ 4.1 Hierarchical Bayesian modelling
Hierarchical models with partial pooling are particularly useful for PBSHM. Because parameters are allowed to vary at the domain level (as opposed to complete pooling), this approach can represent benign variations within a population. In addition, population-level variables are informed by the full dataset, rather than data from a single domain. This increase in statistical power is especially important in situations where one or more domains have limited data A. Gelman et al. (2013); L.A. Bull et al. (2022, 2023a). In such cases, parameters from the data-poor domains exhibit shrinkage towards the population mean (therefore borrowing information from the other domains), which tightens the parameter variance A. Gelman et al. (2013); L.A. Bull et al. (2022, 2023a). The differences among the data-pooling techniques are shown graphically in Figure 4.1, using a simple linear regression example.
@@
In other words, for each observation $i$, the output is determined by evaluating one of $K$ latent functions, $f_{k}\left(x_{i,k}\right)$, plus additive noise, $\epsilon_{i,k}$ L.A. Bull et al. (2023a). While each of the $k$ groups can be learned independently, a combined inference can be used to take advantage of the full $\sum_{k=1}^{K}N_{k}$ population dataset. For example, consider a population that can be expressed using $K$ linear regression models, as in Figure 4.1,
@@
A shared hierarchy of prior distributions can be placed over the slopes and intercepts for the groups $k\in\left\{1,...\,,K\right\}$, in line with a Bayesian framework. To allow information to flow between groups, parent nodes $\left\{\boldsymbol{\mu}_{\alpha},\boldsymbol{\sigma}^{2}_{\alpha}\right\}$ and $\left\{\boldsymbol{\mu}_{\beta},\boldsymbol{\sigma}^{2}_{\beta}\right\}$ can be learned at the population level. Note that sometimes, it may be appropriate to learn certain parameters at the population level rather than the domain level. If, for example, the same hardware was used for data collection within the population, one could assume a global noise variance $\sigma^{2}$ that is the same for each domain. Consider the directed graphical models (DGMs) shown in Figures 1a and 1b. Figure 1a shows the DGM for the linear regression example, given an independent (no-pooling) approach. Each model is learnt independently and no information is shared. In contrast, Figure 1b shows the DGM given a partial pooling approach. The slopes and intercepts are indexed by $k$, and plate notation is used to show that these nodes are repeated. Shared parent nodes $\left\{\boldsymbol{\mu}_{\alpha},\boldsymbol{\sigma}^{2}_{\alpha}\right\}$ and $\left\{\boldsymbol{\mu}_{\beta},\boldsymbol{\sigma}^{2}_{\beta}\right\}$ are outside of the plates and not indexed by $k$, meaning that these are population-level variables, and information is permitted to flow between these and the domain-specific parameters. The noise variance $\sigma^{2}$ is not indexed by $k$, and is also learnt globally, at the population level.
{tikzpicture}
[…]
{tikzpicture}
[…]
@@ 4.2 Modal analysis and FRF estimation
$\textrm{imag}\left[\mathbf{H}_{k}\left(\boldsymbol{\omega}_{k}\right)\right]=-\boldsymbol{\omega}^{2}_{k}\sum_{m=1}^{M}\frac{2A_{m}i\zeta_{(k,m)}\boldsymbol{\omega}_{k}\omega_{nat}^{{(k,m)}}}{\omega_{nat}^{4^{(k,m)}}+\boldsymbol{\omega}^{4}_{k}+2\boldsymbol{\omega}^{2}_{k}\omega_{nat}^{2^{(k,m)}}\left(2\zeta_{(k,m)}^{2}-1\right)}$ (14)
Note that residues $A_{m}$ are not indexed by $k$ in Eq. (13). For the models presented in the case studies below, mode shapes were shared among the population, to address identifiability concerns during sampling, and because mode shapes are often less sensitive to global variations compared to other modal parameters. The current work has focussed on the real part of the FRF, and Eq. (13) provided the mean of the likelihood function for the models developed. The residues, modal damping, and natural frequencies for each domain were then learnt as hyperparameters for the hierarchical model. Learning the real part of the FRF was sufficient for demonstrating the proposed technology. However, the imaginary part could be learnt using the same methods, with Eq. (14) providing the likelihood function, or could be inferred (at least in part), by exploiting the causal relationship between the real and imaginary components of the FRF Worden and G.R. Tomlinson (2001).
@@ 5.1 Data collection at ambient laboratory temperature
The first case used data collected at ambient laboratory temperature, with the blades in a fixed-free boundary condition, which was approximated by placing the root end of each blade in a substantiated strong-wall mount. Ten uniaxial 100 mV/g accelerometers were placed along the length of the underside of each blade. Note that the same accelerometers were used on each blade, and care was taken to ensure that they were attached to approximately the same locations on each blade. An electrodynamic shaker with force gauge was mounted to a fixture bolted to the laboratory floor and attached to the blade 0.575 metres from the root. The shaker was attached to the underside of the blade in the flapwise direction. A continuous random excitation was generated in LMS (note: LMS refers to Siemens PLM LMS SCADAS hardware and software) and applied to excite the blade up to 400 Hz, with a step size of 4.88e-02 Hz. Throughput time data were collected for each test, and the data were divided into 20 blocks. Hanning windows were applied and FRFs were computed in LMS, which were then averaged in the frequency domain. The experimental setup is shown in Figure 5.1, and the accelerometer positions on the blades are shown in Figure 5.1. To simplify the analysis, a narrow frequency band was selected between 24 and 61 Hz, with the fourth and fifth bending modes of the blades dominating the response in this band. Although other modes appear to have a small influence in this band, a 2DOF assumption was imposed. (This assumption results in smoothing of the FRF over the band, and might result in some loss of interpretability, but is acceptable for these preliminary analyses). The real part was modelled as a probabilistic FRF, using the FRF estimate from Eq. (13) as the mean of the likelihood function, as described in Case 1, presented in Section 6 of this paper. The real parts of the averaged FRFs for each blade, at the second accelerometer from the blade root (corresponding to the drive-point location), are shown in Figures 2a and 2b. Figure 2a shows the full measured bandwidth, and Figure 2b shows the FRF in the bandwidth of interest, between 24 and 61 Hz.
@@
Figures 2a and 2b show increasing variability with respect to frequency, which is an expected result, given that higher-frequency modes are more sensitive to small physical changes than lower-frequency modes. For modes less than 80 Hz, the maximum frequency difference among the blades was approximately 2.5 Hz; for modes greater than 80 Hz, the maximum frequency difference was approximately 6.3 Hz. Note the grouping visible at several of the peaks, where Blades 1 and 2 appear closely aligned in frequency while Blades 3 and 4 appear closely aligned. These results are quite relevant for PBSHM. All of the helicopter blades are healthy, and represent a normal-condition state of the population. Consider a situation where only FRFs from one of the groupings are available for training a model (or FRFs from the other groups are missing data). The normal condition could be heavily biased towards the training set, and incoming FRFs could be flagged as damaged, even if they are healthy. Further details regarding the data collection and processing for these tests can be found in T.A. Dardeno et al. (2022).
@@ 5.2 Data collection at various temperatures in an environmental chamber
The second case used data from Blade 1, collected at temperatures ranging from -20 to ${}^{\circ}$ in increments of ${}^{\circ}$ in an environmental chamber. The blade was tested in an approximately fixed-free boundary condition, with the blade root mounted on a fixture that was substantiated with a large concrete block. These tests used the same accelerometers and sensor layout, as the previous tests at ambient laboratory temperature. Likewise, the same data-acquisition parameters and processing methods were used. The shaker was mounted to the bottom plate of the test fixture, and the shaker and force gauge were connected to the underside of the blades to excite the blades in the flapwise direction. A thermal jacket was used to protect the shaker when testing at lower temperatures. Prior to data collection, the environmental chamber was set to each of the temperatures of interest, after which the blades were allowed to soak for at least two hours to reach the desired temperature. Throughput force and acceleration data were collected for each test. The data were then divided into 20 blocks, and a Hanning window was applied to each data block. FRFs were then computed in LMS, and averaged in the frequency domain. The experimental setup, with the helicopter blade inside the environmental chamber, is shown in Figure 3. To simplify the analysis, a narrow frequency band was selected between 135 and 155 Hz, with a higher-order bending mode of the blade dominating the response in this band. An SDOF assumption was imposed. (Again, this assumption was considered acceptable for these preliminary analyses, to avoid data-separability issues.) FRFs captured at temperatures -10, -5, 10, and ${}^{\circ}$ provided the training data. The remaining temperature-varied FRFs were used to evaluate the extrapolation results to other temperatures. As with the previous case, the real part was modelled as a probabilistic FRF, using the FRF estimate from Eq. (13) as the mean of the likelihood function. The real parts of the averaged FRFs for each temperature, from the fourth accelerometer from the blade tip, are shown in Figures 4a and 4b. Figure 4a shows the full measured bandwidth, and Figure 4b shows the FRF in the bandwidth of interest, between 135 and 155 Hz.
@@ 6 Case 1: population-based modelling of FRFs for nominally-identical structures
The first case used FRF data from four nominally-identical helicopter blades, collected at ambient laboratory temperature. A population form for the FRFs of the helicopter blades was developed using the aforementioned hierarchical (partial-pooling) approach. Models were developed using the probabilistic programming language Stan. Analyses were performed using MCMC, via the no U-turn (NUTS) implementation of Hamiltonian Monte Carlo (HMC) M.D. Hoffman and Gelman (2014); M. Betancourt and Girolami (2015). HMC uses approaches based in differential geometry to generate transitions that span the full marginal variance, which allows the algorithm to accommodate large curvature changes in the target distribution (which are common with hierarchical models) M. Betancourt and Girolami (2015). This allows the sampler to efficiently explore the posterior distribution, without being susceptible to the random-walk behaviour that can occur with other samplers M. Betancourt and Girolami (2015).
@@ 6.1 Model development
where for each observation $i$, the output was determined by evaluating one of $K$ latent functions, $f_{k}\left(\omega_{i,k}\right)$, plus additive noise, $\epsilon_{i,k}$. For this work, there were four helicopter blades and four domains. Therefore, the population model included four latent functions (i.e., $K=4$). The (real) FRFs were modelled probabilistically with an assumed Gaussian-distributed likelihood,
$\text{real}\left[\mathbf{H}_{k}\left(\boldsymbol{\omega}_{k}\right)\right]\sim\mathcal{N}\left(\mathbf{f}_{k}\left(\boldsymbol{\omega}_{k}\right),\sigma^{2}_{\text{H}}\right)$ (18)
where $\mathbf{f}_{k}\left(\boldsymbol{\omega}_{k}\right)$ was equal to the real component of the FRF, calculated using modal parameters via Eq. (13). The additive noise variance $\sigma^{2}_{\text{H}}$ of the FRFs was assumed to be the same for each blade. This assumption was reasonable, as the same data acquisition system and sensors were used among the different tests. Note that $\boldsymbol{\omega}_{k}$ was permitted to vary depending on the given FRF. This allowed for flexibility, to consider FRFs with different numbers of spectral lines or missing data points, and to represent population uncertainty. Natural frequencies and modal damping were learnt at the domain level, and allowed to vary among the different helicopter blades. Residues were shared among the different domains and learnt at the population level, to mitigate model identifiability issues. Shared, population-level prior distributions (with hyperpriors) were also placed over the modal parameters to capture/infer the similarity among the FRFs. Domain-level natural frequencies, $\boldsymbol{\omega}_{nat}=\{\{\omega_{nat}^{k,m}\}_{m=1}^{2}\}_{k=1}^{4}$, were sampled from a truncated normal parent distribution, with higher-level expectation and variance sampled from truncated normal distributions,
@@
Domain-level damping, $\boldsymbol{\zeta}=\{\{\zeta_{k,m}\}_{m=1}^{2}\}_{k=1}^{4}$, was sampled from a beta parent distribution, with higher-level shape parameters sampled from truncated normal distributions,
$\begin{split}&\boldsymbol{\zeta}\sim\mathcal{B}\left(\boldsymbol{\alpha}_{\zeta},\boldsymbol{\beta}_{\zeta}\right)\\ &\boldsymbol{\alpha}_{\zeta}\sim\mathcal{TN}\left([6,6],[0.5^{2},0.5^{2}]\right)\\ &\boldsymbol{\beta}_{\zeta}\sim\mathcal{TN}\left([1000,1000],[10^{2},10^{2}]\right)\\ \end{split}$ (20)
@@
$\begin{split}&\boldsymbol{A}\sim\mathcal{TN}\left(\boldsymbol{\mu}_{A},\boldsymbol{\sigma}^{2}_{A}\right)\\ &\boldsymbol{\mu}_{A}\sim\mathcal{N}\left([\text{-}0.004,\text{-}0.004],[0.003^{2},0.003^{2}]\right)\\ &\boldsymbol{\sigma}_{A}^{2}\sim\mathcal{TN}\left([0.003,0.003],[0.003^{2},0.003^{2}]\right)\\ \end{split}$ (21)
For simplicity, normal distributions were chosen for most of the parameters (or truncated normal distributions, if the parameter was restricted to be positive), although other distributions could be used. A beta distribution was chosen for damping, as beta distributions have support $x\in[0,1]$, which is suitable for lightly-damped systems. Hyperpriors were chosen based on the physics that could be interpreted by visual inspection of the training data (e.g., natural frequency), or by fitting the data-rich domains independently (discussed below). Note that hyperpriors for $\boldsymbol{\omega}_{nat}$ are shown in rad/s. Shared noise variance, $\sigma^{2}_{\text{H}}$, was sampled from a truncated normal parent distribution, with higher-level expectation and variance also sampled from truncated normal distributions,
$\begin{split}&\sigma_{H}{{}^{2}}\sim\mathcal{TN}\left(\mu_{\sigma^{2}},\sigma^{2}_{\sigma^{2}}\right)\\ &\mu_{\sigma^{2}}\sim\mathcal{TN}\left(0,100^{2}\right)\\ &\sigma^{2}_{\sigma^{2}}\sim\mathcal{TN}\left(100,100^{2}\right)\\ \end{split}$ (22)
Note that a non-informative prior was used for the noise variance, to stabilise computation. To incorporate more prior information, the half-t family of distributions (e.g., half-Cauchy distribution) can be used instead A. Gelman (2006). A graphical model displaying the parameter hierarchy is shown in Figure 5.
{tikzpicture}
[…]

### v2

@@ 3 Novelty and Contribution
Unlike previous efforts, the current work is concerned with developing a probabilistic FRF model, using an FRF equation based on modal parameters (natural frequency, mode shape, and damping) as the likelihood function. As with L.A. Bull et al. (2022, 2023a, 2023b), a hierarchical Bayesian (or MTL) framework was used. However, the focus is data sharing in the presence of low-resolution frequency information (i.e., given sparse data). The inability to localise both time and frequency to a fine resolution, per the uncertainty principle, means that missing time-domain data (e.g., acceleration, velocity) measured from SHM sensors will result in fewer spectral lines in the frequency domain. This decreased frequency resolution could impair proper characterisation/identification of modal peaks, which may be features of interest in a damage-detection application. By using a combined-inference approach, structures whose FRFs have many spectral lines in a band of interest can lend statistical support to those whose FRFs are limited by missing data. In addition, the current work incorporated functional relationships to describe changes in environmental conditions. This inclusion of functional relationships allows for extrapolation to temperature states not used in model training, which increases the amount of normal-condition information available, to address data-scarcity challenges. In contrast to L.A. Bull et al. (2023b), parameters were learnt over the experimental campaign, rather than hyperparameters, giving greater physical interpretability of the results. In contrast to previous work, the current work focusses on information transfer for modal analysis and meta-analysis prediction for future models, rather than data pooling alone. Beyond conventional multilevel modelling and partial pooling, the intent is to learn how parameters vary over the population, even in the absence of complete data - an area that has not yet been addressed in modal analysis. Again, the current work allows for the prediction of new models, in addition to unseen data.
@@ 4 Background theory
In this work, combined probabilistic FRF models were developed using a hierarchical Bayesian modelling approach, to support SHM research. This section provides an overview of the methodologies used, including hierarchical multilevel modelling from a Bayesian perspective, linear modal analysis, FRF estimation, and model evaluation.
@@ 4.1 Hierarchical Bayesian modelling
Hierarchical models with partial pooling can be applied in a variety of SHM contexts, where a system can benefit from information sharing while preserving the uniqueness of the individual data sources. However, these models are particularly useful for PBSHM, and link intuitively to the concept of the population form, which encompasses both the essential nature and variabilities among the members of a homogeneous population. Because parameters are allowed to vary at the domain level (as opposed to complete pooling), this approach can represent benign variations within a population. In addition, population-level variables are informed by the full dataset, rather than data from a single domain. This increase in statistical power is especially important in situations where one or more domains have limited data A. Gelman et al. (2013); L.A. Bull et al. (2022, 2023a). In such cases, parameters from the data-poor domains exhibit shrinkage towards the population mean (therefore borrowing information from the other domains), which tightens the parameter variance A. Gelman et al. (2013); L.A. Bull et al. (2022, 2023a). The differences among the data-pooling techniques are shown graphically in Figure 4.1, using a simple linear regression example.
@@
In other words, for each $i^{th}$ observation, the output is determined by evaluating one of $K$ latent functions, $f_{k}\left(x_{i,k}\right)$, plus additive noise, $\epsilon_{i,k}$ L.A. Bull et al. (2023a). While each of the $k$ groups can be learned independently, a combined inference can be used to take advantage of the full $\sum_{k=1}^{K}N_{k}$ population dataset. For example, consider a population that can be expressed using $K$ linear regression models, as in Figure 4.1,
@@
A shared hierarchy of prior distributions can be placed over the slopes and intercepts for the groups $k\in\left\{1,...\,,K\right\}$, in line with a Bayesian framework. To allow information to flow between groups, parent nodes $\left\{\boldsymbol{\mu}_{\alpha},\boldsymbol{\sigma}^{2}_{\alpha}\right\}$ and $\left\{\boldsymbol{\mu}_{\beta},\boldsymbol{\sigma}^{2}_{\beta}\right\}$ can be learned at the population level. Note that sometimes, it may be appropriate to learn certain parameters at the population level rather than the domain level. If, for example, the same hardware was used for data collection within the population, one could assume a global noise variance ($\sigma^{2}$) that is the same for each domain. Consider the directed graphical models (DGMs) shown in Figures 1a and 1b. Figure 1a shows the DGM for the linear regression example, given an independent (no-pooling) approach. Each model is learnt independently and no information is shared. Slopes, intercepts, and noise variance are learnt individually for each $k$ domain, but there is no plate surrounding them to denote repetition over the $K$ domains. While the slopes and intercepts are sampled from shared parent nodes $\left\{\boldsymbol{\mu}_{\alpha},\boldsymbol{\sigma}^{2}_{\alpha}\right\}$ and $\left\{\boldsymbol{\mu}_{\beta},\boldsymbol{\sigma}^{2}_{\beta}\right\}$, these nodes are informed only by the data from a single domain and any priors. In contrast, Figure 1b shows the DGM given a partial-pooling approach. The slopes and intercepts are indexed by $k$, and plate notation is used to show that these nodes are repeated. Once again, shared parent nodes are outside of the plates and not indexed by $k$, meaning that these are population-level variables, and information is permitted to flow between these and the domain-specific parameters. In other words, the parent nodes are now informed by the data from multiple domains and prior information. The noise variance ($\sigma^{2}$) is shared among the domains, at the population level, as indicated by the absence of a $k$ subscript, and its exclusion from the $K$ plate.
$y_{i,k}$$x_{i,k}$$\boldsymbol{\alpha}_{k}$$\boldsymbol{\beta}_{k}$$\mu_{\alpha}$$\sigma^{2}_{\alpha}$$\mu_{\beta}$$\sigma^{2}_{\beta}$$\sigma_{k}^{2}$$i\in 1:N_{k}$
$y_{i,k}$$x_{i,k}$$\boldsymbol{\alpha}_{k}$$\boldsymbol{\beta}_{k}$$\mu_{\alpha}$$\sigma^{2}_{\alpha}$$\mu_{\beta}$$\sigma^{2}_{\beta}$$\sigma^{2}$$i\in 1:N_{k}$$k\in 1:K$
@@ 4.2 Modal analysis and FRF estimation
$\textrm{imag}\left[\mathbf{H}_{k}\left(\boldsymbol{\omega}_{k}\right)\right]=-\boldsymbol{\omega}^{2}_{k}\sum_{m=1}^{M}\frac{-2A_{m}i\zeta_{(k,m)}\boldsymbol{\omega}_{k}\omega_{nat}^{{(k,m)}}}{\omega_{nat}^{4^{(k,m)}}+\boldsymbol{\omega}^{4}_{k}+2\boldsymbol{\omega}^{2}_{k}\omega_{nat}^{2^{(k,m)}}\left(2\zeta_{(k,m)}^{2}-1\right)}$ (14)
Note that residues, $A_{m}$, are not indexed by $k$ in Eq. (13). For the models presented in the case studies below, mode shapes were shared among the domains, to address identifiability concerns during sampling, and because mode shapes are often less sensitive to global variations compared to other modal parameters. The current work has focussed on the real part of the FRF, and Eq. (13) provided the mean of the likelihood function for the models developed. The residues, modal damping, and natural frequencies for each domain were learnt as parameters for the hierarchical model, with distributions over these parameters learnt as hyperparameters. Learning the real part of the FRF was sufficient for demonstrating the proposed technology. However, the imaginary part could be learnt using the same methods, with Eq. (14) providing the likelihood function, or could be inferred (at least in part), by exploiting the causal relationship between the real and imaginary components of the FRF Worden and G.R. Tomlinson (2001).
@@ 5.1 Data collection at ambient laboratory temperature
The first case used data collected at ambient laboratory temperature, with the blades in a fixed-free boundary condition, which was approximated by placing the root end of each blade in a substantiated strong-wall mount. Ten uniaxial 100 mV/g accelerometers were placed along the length of the underside of each blade. Note that the same accelerometers were used on each blade, and care was taken to ensure that they were attached to approximately the same locations on each blade. An electrodynamic shaker with force gauge was mounted to a fixture bolted to the laboratory floor and attached to the blade 0.575 metres from the root. The shaker was attached to the underside of the blade in the flapwise direction. A continuous random excitation was generated in LMS (note: LMS refers to Siemens PLM LMS SCADAS hardware and software) and applied to excite the blade up to 400 Hz, with a step size of 4.88e-02 Hz. Throughput time data were collected for each test, and the data were divided into 20 blocks. Hanning windows were applied and FRFs were computed for each time block, which were then averaged in the frequency domain. The experimental setup is shown in Figure 5.1, and the accelerometer positions on the blades are shown in Figure 5.1. To demonstrate the proposed technique and reduce the number of model hyperparameters (thus minimising model complexity and reducing computation time), a narrow frequency band was selected around two lightly-damped, well-separated modes. Specifically, the chosen bandwidth was between 24 and 61 Hz, with the fourth and fifth bending modes of the blades dominating the response in this band. Although other modes appear to have a small influence in this band, a 2DOF assumption was imposed. (This assumption results in smoothing of the FRF over the band, and might result in some loss of interpretability, but is acceptable for these preliminary analyses). The real part was modelled as a probabilistic FRF, using the FRF estimate from Eq. (13) as the mean of the likelihood function, as described in Case 1, presented in Section 6 of this paper. The real parts of the averaged FRFs for each blade, at the second accelerometer from the blade root (corresponding to the drive-point location), are shown in Figures 2a and 2b. Figure 2a shows the full measured bandwidth, and Figure 2b shows the FRF in the bandwidth of interest, between 24 and 61 Hz.
@@
Figures 2a and 2b show increasing variability with respect to frequency, which is an expected result, as higher-frequency modes are more sensitive to small physical changes than lower-frequency modes. For modes less than 80 Hz, the maximum frequency difference among the blades was approximately 2.5 Hz; for modes greater than 80 Hz, the maximum frequency difference was approximately 6.3 Hz. Note the grouping visible at several of the peaks, where Blades 1 and 2 appear closely aligned in frequency while Blades 3 and 4 appear closely aligned. These results are quite relevant for PBSHM. All of the helicopter blades are healthy, and represent a normal-condition state of the population. Consider a situation where only FRFs from one of the groupings are available for training a model (or FRFs from the other groups are missing data). The normal condition could be heavily biased towards the training set, and incoming FRFs could be flagged as damaged, even if they are healthy. Further details regarding the data collection and processing for these tests can be found in T.A. Dardeno et al. (2022).
@@ 5.2 Data collection at various temperatures in an environmental chamber
The second case used data from Blade 1, collected at temperatures ranging from -20 to ${}^{\circ}$ in increments of ${}^{\circ}$ in an environmental chamber. The blade was tested in an approximately fixed-free boundary condition, with the blade root mounted on a fixture that was substantiated with a large concrete block. These tests used the same accelerometers, sensor layout, data-acquisition parameters, and processing methods as the previous tests at ambient laboratory temperature. The shaker was mounted to the bottom plate of the test fixture, and the shaker and force gauge were connected to the underside of the blades to excite the blades in the flapwise direction. A thermal jacket was used to protect the shaker when testing at lower temperatures. Prior to data collection, the environmental chamber was set to each of the temperatures of interest, after which the blades were allowed to soak for at least two hours to reach the desired temperature. Throughput force and acceleration data were collected for each test. The data were then divided into 20 blocks, and a Hanning window was applied to each data block. FRFs were then computed for each block, and averaged in the frequency domain. The experimental setup, with the helicopter blade inside the environmental chamber, is shown in Figure 3. To minimise the number of hyperparameters in the model and reduce computation time, a narrow frequency band was selected around a lightly-damped, well-separated modal peak. Specifically, the chosen bandwidth was between 135 and 155 Hz, with a higher-order bending mode dominating the response in this band. An SDOF assumption was imposed. (Again, this assumption was considered acceptable for these preliminary analyses, to avoid data-separability issues. Likewise, the coefficients of the functions that relate temperature fluctuations to changes in modal parameters will likely be mode-dependent.) FRFs captured at temperatures -10, -5, 10, and ${}^{\circ}$ provided the training data. The remaining temperature-varied FRFs were used to evaluate the extrapolation results to other temperatures. As with the previous case, the real part was modelled as a probabilistic FRF, using the FRF estimate from Eq. (13) as the mean of the likelihood function. The real parts of the averaged FRFs for each temperature, from the fourth accelerometer from the blade tip, are shown in Figures 4a and 4b. Figure 4a shows the full measured bandwidth, and Figure 4b shows the FRF in the bandwidth of interest, between 135 and 155 Hz.
@@ 6 Case 1: population-based modelling of FRFs for nominally-identical structures
The first case used FRF data from four nominally-identical helicopter blades, collected at ambient laboratory temperature. A population form for the FRFs of the helicopter blades was developed using the aforementioned hierarchical (partial-pooling) approach. Models were developed using the probabilistic programming language Stan. Analyses were performed using MCMC, via the no U-turn (NUTS) implementation of Hamiltonian Monte Carlo (HMC) M.D. Hoffman and Gelman (2014); M. Betancourt and Girolami (2015). HMC uses approaches based in differential geometry to generate transitions that span the full marginal variance, which allows the algorithm to accommodate large curvature changes in the target distribution (which are common with hierarchical models) M. Betancourt and Girolami (2015).
@@ 6.1 Model development
where for each $i^{th}$ observation, the output was determined by evaluating one of $K$ latent functions, $f_{k}\left(\omega_{i,k}\right)$, plus additive noise, $\epsilon_{i,k}$. For this work, there were four helicopter blades and four domains. Therefore, the population model included four latent functions (i.e., $K=4$). To accommodate outliers in the FRF data, the (real) FRFs were modelled probabilistically with an assumed Student’s t-distributed likelihood,
$\text{real}\left[\mathbf{H}_{k}\left(\boldsymbol{\omega}_{k}\right)\right]\sim\mathcal{T}\left(\nu,\mathbf{f}_{k}\left(\boldsymbol{\omega}_{k}\right),\sigma^{2}\right)$ (18)
where $\mathbf{f}_{k}\left(\boldsymbol{\omega}_{k}\right)$ was equal to the real component of the FRF, calculated using modal parameters via Eq. (13). For both the independent and partial-pooling models, the degrees-of-freedom parameter, $\nu$, was assumed to be four (4). Setting $\nu$ to a constant was necessary to stabilise the independent (no-pooling) models for the data-poor domains, to avoid the sampler falsely identifying the modal peaks as noise. (Note that struggling to fit the $\nu$ parameter was not a problem for the data-rich independent models or for the partial-pooling model, only for the data-poor independent models.) The additive noise variance, $\sigma^{2}$, of the FRFs was assumed to be the same for each blade. This assumption was reasonable, as the same data acquisition system and sensors were used among the different tests. Note that $\boldsymbol{\omega}_{k}$ was permitted to vary depending on the given FRF. This allowed for flexibility, to consider FRFs with different numbers of spectral lines or missing data points, and to represent population uncertainty. Natural frequencies and modal damping were learnt at the domain level, and allowed to vary among the different helicopter blades. Residues were shared among the different domains and learnt at the population level, to mitigate model identifiability issues. Shared, population-level prior distributions (with hyperpriors) were also placed over the modal parameters to capture/infer the similarity among the FRFs. Domain-level natural frequencies, $\boldsymbol{\omega}_{nat}=\{\{\omega_{nat}^{k,m}\}_{m=1}^{2}\}_{k=1}^{4}$, were sampled from a truncated normal parent distribution, with higher-level expectation and variance sampled from truncated normal distributions,
@@
Note that hyperpriors for $\boldsymbol{\omega}_{nat}$ are shown in rad/s. Domain-level damping, $\boldsymbol{\zeta}=\{\{\zeta_{k,m}\}_{m=1}^{2}\}_{k=1}^{4}$, was sampled from a beta parent distribution, with higher-level shape parameters sampled from truncated normal distributions,
$\begin{split}&\boldsymbol{\zeta}\sim\mathcal{B}\left(\boldsymbol{\alpha}_{\zeta},\boldsymbol{\beta}_{\zeta}\right)\\ &\boldsymbol{\alpha}_{\zeta}\sim\mathcal{TN}\left([6,6],[0.5^{2},0.5^{2}]\right)\\ &\boldsymbol{\beta}_{\zeta}\sim\mathcal{TN}\left([500,500],[50^{2},50^{2}]\right)\\ \end{split}$ (20)
@@
$\begin{split}&\boldsymbol{A}\sim\mathcal{TN}\left(\boldsymbol{\mu}_{A},\boldsymbol{\sigma}^{2}_{A}\right)\\ &\boldsymbol{\mu}_{A}\sim\mathcal{N}\left([\text{-}0.004,\text{-}0.004],[0.0035^{2},0.0035^{2}]\right)\\ &\boldsymbol{\sigma}_{A}^{2}\sim\mathcal{TN}\left([0.003,0.003],[0.003^{2},0.003^{2}]\right)\\ \end{split}$ (21)
For simplicity, normal distributions were chosen for most of the parameters (or truncated normal distributions, if the parameter was restricted to be positive), although other distributions could be used. A beta distribution was chosen for damping, as beta distributions have support $x\in[0,1]$, which is suitable for lightly-damped systems. Hyperpriors were chosen based on the physics that could be interpreted by visual inspection of the training data (e.g., natural frequency), or by fitting the data-rich domains independently (discussed below). Shared noise variance, $\sigma^{2}$, was sampled from a half-Cauchy parent distribution, with higher-level scale parameter, $\gamma^{2}$, sampled from a truncated normal distribution,
$\begin{split}&\sigma^{2}\sim\mathcal{C}\left(0,\gamma^{2}\right)\\ &\gamma^{2}\sim\mathcal{TN}\left(0.01,0.05^{2}\right)\\ \end{split}$ (22)
The half-Cauchy was considered an appropriate prior distribution for the shared noise term, as it peaks at zero, but has a long right tail, to allow for some flexibility in the assumptions A. Gelman (2006). (Note that, in practice, measurement noise would be expected to be fairly small given that FRFs are typically averaged.) A graphical model displaying the parameter hierarchy is shown in Figure 5.
$H_{i,k}$$\omega_{i,k}$$\boldsymbol{\zeta}_{k}$$\boldsymbol{\omega}_{nat}^{k}$$\boldsymbol{A}$$\boldsymbol{\mu}_{\omega_{nat}}$$\boldsymbol{\sigma}^{2}_{\omega_{nat}}$$\boldsymbol{\alpha}_{\zeta}$$\boldsymbol{\beta}_{\zeta}$$\boldsymbol{\mu}_{A}$$\boldsymbol{\sigma}^{2}_{A}$$\sigma^{2}$$\gamma^{2}$ $\nu\text{ }$$i\in 1:N_{k}$$k\in 1:K$